OpenStudy (anonymous):

How would the expression x^3+27 be rewritten using Sum of Cubes?

OpenStudy (vishweshshrimali5):

The basic formula is this: \[\large{a^3 + b^3 = (a+b)(a^2-ab+b^2)}\]

OpenStudy (vishweshshrimali5):

Now your expression can be written as: \[\large{x^3 + 27 = x^3 + 3^3}\] Compare this with the formula I mentioned above

OpenStudy (anonymous):

but i dont understand because there isnt an answer like that. A.(x+3)(x^2+3x-9) B.(x+3)(x^2-3x+9) C.(x-3)(x^2-3x+9) D.(x+3)(x^2-3x-9)